OpenStudy (anonymous):

How do I reach Pi again? I kinda forgot and now I have a serious problem. <_> Help please.

OpenStudy (anonymous):

What is the problem?

OpenStudy (anonymous):

xD Of course you would answer first that's very nice, actually I'm just trying to remember because my teacher is going to have us work on cylindrical volumes again. just a review... I don't wanna be behind in class like that.

OpenStudy (anonymous):

Oh, okay, I wont answer you anymore then.

OpenStudy (anonymous):

heh, don't joke like that.

OpenStudy (anonymous):

Im serious, Why do you think im joking?

OpenStudy (anonymous):

lol cause I didn't quite say anything offensive... I don't think I did at least...

OpenStudy (anonymous):

You said "of course you would answer first..."

OpenStudy (anonymous):

xD Of course you would answer first that's very nice... that's very nice... nice

OpenStudy (anonymous):

:l

OpenStudy (anonymous):

Oh ok, I thought you were being sarcastic :/ ..but im sleepy. I'm signing off, going to bed, goodnight @Guyofreckoning

OpenStudy (anonymous):

Night. Take it easy.

OpenStudy (anonymous):

Well... What is the problem? :)

OpenStudy (anonymous):

just needa remember the equation to get to Pi...

OpenStudy (anonymous):

You want to approximate \(\pi\)?

OpenStudy (anonymous):

yeah. something with division and circles, can't remember the thing for the life of me.

OpenStudy (anonymous):

Well pi is the ration between the circumference (perimeter of circle) and the diameter.

OpenStudy (anonymous):

ok I think I can do this now, dumb old me forgot about google. heh... but thanks anyways!

OpenStudy (richyw):

This converges to \(\pi\). Start now and in your lifetime you might get it to like 5 decimal places :P\[\pi= 4 \sum_{n=0}^{\infty} \frac{2}{(4n+1)(4n+3)}\]Seriously though \(\pi\) is just \(\pi\). It should always be kept as that unless you want a decimal approximation of something. in that case \(\pi\) is just what you can memorize. You can't compute the decimal value of pi because it is irrational.

OpenStudy (anonymous):

O: I dunno if I should be so mean as to switch the best result. but that is one heck of a reply

OpenStudy (anonymous):

sry wio :c but that is the best reply.